OpenStudy (anonymous):

*Product of a Polynomial* Find the product: (3a-2b)(3a+3b)

OpenStudy (anonymous):

(3a−2b)(3a+3b) =(3a+−2b)(3a+3b) =(3a)(3a)+(3a)(3b)+(−2b)(3a)+(−2b)(3b) =9a2+9ab−6ab−6b2 =9a2+3ab−6b2

OpenStudy (anonymous):

i just simplified it

OpenStudy (anonymous):

Can you help me with all of these? There are only ten questions and I honestly cannot understand these, these are the only ones I'm having problems with. :( (I'm homeschooled)

OpenStudy (bee_see):

you have to use FOIL for these problems